Find a partial differential equation by eliminating a and b from the equations

z^2 = ax^3 + by^3 + ab.


1
Expert's answer
2022-03-07T04:52:01-0500

Let us find a partial differential equation by eliminating "a" and "b" from the equations "z^2 = ax^3 + by^3 + ab."

Let us use the implicit differentiation.

It follows that

"2zz_x = 3ax^2" and "2zz_y = 3by^2."

Therefore, "a=\frac{2zz_x}{ 3x^2}" and "b=\frac{2zz_y}{ 3y^2}."

We conclude that the partial differential equation is of the form:

"z^2 = \frac{2zz_x}{ 3x^2}x^3 + \frac{2zz_y}{ 3y^2}y^3 + \frac{2zz_x}{ 3x^2}\frac{2zz_y}{ 3y^2}."